OpenStudy (anonymous):

Simplify this. 5(x - 2) + 6 =

OpenStudy (anonymous):

5x - 10 + 6 = 5x - 4 =

OpenStudy (anonymous):

thanks but could you explain to me how you got that answer?

OpenStudy (anonymous):

5(x-2) + 6 = 0 5x - 10 + 6 = 0 5x + 6 = 10 5x = 4 x = 4/5

OpenStudy (anonymous):

The above is correct if that is all equal to zero. I didn't assume it was because you didn't say = 0 you just said =. You distribute the 5 by multiplying it by each item in parenthesis.